OpenStudy (anonymous):

Which of the following graphs correctly represents the inequality y<- (-3/4x)+2 ?

OpenStudy (anonymous):

is that \[ \frac{ 3x }{ 4 } or \frac{ 3 }{ 4x }\]

OpenStudy (anonymous):

\[y \le -\frac{ 3 }{ 4 }x+2\]

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

u want the answer or the explanation?

OpenStudy (anonymous):

is it the third graph?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

ty